# A Quality Diversity Method to Automatically Generate Multi-Agent Path Finding Benchmark Maps

### v3

@@ A.1 General Hardness
Number of Obstacles vs. Std of BC. We run the one-algorithm experiment of PBS with the number of obstacles and std of BC as the diversity measures to explore the relationship between the two measures and their correlation with the general hardness of generated maps. From Figure 10, we observe that in the y-axis, with the increase of obstacles, the CPU runtime of PBS is increasing, which is consistent with the one-algorithm experiment result of PBS in Section 4 of the main text. With fewer variations in the CPU runtime of PBS at a given number of obstacles, we can quantify the general hardness of the maps based on the number of obstacles. From the x-axis, however, there are many maps with the same std of BC but showing different general hardness of the maps for PBS. Therefore, the std of BC is not a good measure to quantify the general hardness of the maps.
@@ A.2 Spatial Arrangement
KL Divergence of Tile Pattern Distribution vs. Map Entropy. We run a one-algorithm experiment with PBS. We use the same objective in Section 4 and use map entropy and KL divergence of the tile pattern distribution as the diversity measures. Figure 12 shows the resulting archive of maps. We observe that, with map entropy less than 0.5, the generated maps are usually composed of a large chunk of empty spaces with a large component of obstacles on the boundary, revealing a simple structure. Compared to maps generated by the KL divergence of the tile distribution in Figure 5 of Section 4, maps generated by map entropy have less diverse patterns. Therefore, we use the KL divergence of tile pattern distribution as the diversity measure.
@@
KL Divergence ofTile Pattern Distribution vs. KL Divergence of WL Graph Feature. We run the one-algorithm experiment of PBS with diversity measures of the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature to explore their correlation and the spatial arrangements in generated maps. From Figure 13, we observe that all generated maps are well-distributed across the archive, revealing a similar distribution of generated maps with different pairs of KL divergence of the tile pattern distribution and the KL divergence of WL graph feature. Therefore, these two diversity measures are similar. Both quantify the similarity between the generated maps and the pre-defined set of maze maps (Stern et al. 2019). We use the KL divergence of the tile pattern distribution in our experiments since the WL graph feature needs a graph transformation and several iterations of hash transformation, which is more computationally expensive.
Appendix B Maps with similar patterns
Figure 14: Similar maps with evenly distributed short corridor components or large chunks of empty space on which CBS performs well. Maps (a) and (b) are maps with evenly distributed short corridor components. Map (c) is the map with large chunks of empty space. Map (d) is the map with both evenly distributed short corridor components and large chunks of empty space, but also long corridors on the left.
Figure 14 shows the easy maps of CBS, similar to the maps shown in Figure 3 in Section 4 in the main text.
Figure 15: Map (a) and Map (b) are maps similar to representative maps with long corridors and one-tile spaces on which EECBS performs poorly. Map (c) and Map (d) are maps similar to representative maps with two or more columns of empty space between each long obstacle component and short obstacle components on which EECBS performs well.
Figure 15 shows the challenging and easy maps of EECBS, similar to the maps shown in Figure 4 in Section 4 in the main text.
@@ Appendix B Maps with similar patterns
Figure 16 shows the challenging and easy maps of PBS, similar to the maps shown in Figure 5 in Section 4 in the main text.
@@
Figure 17 shows the challenging and easy maps of PIBT, similar to the maps shown in Figure 6 in Section 4 in the main text.
@@
Figure 18 shows the challenging and easy maps of LTF, similar to the maps shown in Figure 7 in Section 4 in the main text.

### v4

@@ A.1 General Hardness
Number of Obstacles vs. Std of BC. To compare the ability to quantify the general hardness of the maps between the number of obstacles and the std of BC, we generate 4370 maps with our pipeline with the objective set as similarity score Zhang et al. [2023a] between repaired and unrepaired maps. We then set the diversity measure as the number of obstacles and std of BC. In this way, we generate a collection of benchmark maps that are not biased towards any algorithm while making sure that they are diverse in terms of the two measures of interest. Then we run PBS on all the generated maps, collecting average CPU runtime over $N_{e}=5$ MAPF instances and the std of BC. We plot them as a scatter plot in Figure 10. We observe that in the y-axis, with the increase of obstacles, the CPU runtime of PBS is increasing, which is consistent with the one-algorithm experiment result of PBS in Section 4 of the main text. With fewer variations in the CPU runtime of PBS at a given number of obstacles, we can quantify the general hardness of the maps based on the number of obstacles. From the x-axis, however, there are many maps with the same std of BC but showing different CPU runtimes of the maps for PBS. Therefore, we argue that the number of obstacles is a better measure than the std of BC to quantify the general hardness of the maps.
@@ A.2 Spatial Arrangement
KL Divergence of Tile Pattern Distribution vs. Map Entropy. We run a one-algorithm experiment with PBS. We use the same objective in Section 4 and use map entropy and KL divergence of the tile pattern distribution as the diversity measures. Figure 12 shows the resulting archive of maps. We observe that, with map entropy less than 0.5, the generated maps are usually composed of a large chunk of empty spaces with a large component of obstacles on the boundary, revealing a simple structure. Compared to maps generated by the KL divergence of the tile distribution in Figure 5 of Section 4, maps generated by map entropy, such as Map (a) and (b), have less diverse patterns. Therefore, we use the KL divergence of tile pattern distribution as the diversity measure.
@@
KL Divergence of Tile Pattern Distribution vs. KL Divergence of WL Graph Feature. We run the one-algorithm experiment of PBS with diversity measures of the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature to explore their correlation and the spatial arrangements in generated maps. From Figure 13 , we observe that all generated maps are well-distributed across the archive, revealing a similar distribution of generated maps with different pairs of KL divergence of the tile pattern distribution and the KL divergence of WL graph feature. Therefore, these two diversity measures are similar. Both quantify the similarity between the generated maps and the pre-defined set of maze maps Stern et al. [2019]. We use the KL divergence of the tile pattern distribution in our experiments since the WL graph feature needs a graph transformation and several iterations of hash transformation, which is more computationally expensive.
Appendix B Implementation and Compute Resource
Implementation. We implement CMA-MAE with Pyribs Tjanaka et al. [2023]. For MAPF algorithms, we use the open source implementations of CBS and EECBS [^3] https://github.com/Jiaoyang-Li/EECBS, PBS [^4] https://github.com/Jiaoyang-Li/PBS, PIBT [^5] https://github.com/Kei18/pibt2, and LTF [^6] https://github.com/AIRI-Institute/learn-to-follow.
Appendix C Maps with similar patterns
Figure 14: Similar maps with evenly distributed short corridor components or large chunks of empty space on which CBS can solve. Maps (a) and (b) are maps with evenly distributed short corridor components. Map (c) and Map (d) are maps with large chunks of empty space on the right, but also long corridors on the left.
Figure 15: Map (a) and Map (b) are maps similar to representative maps with long corridors and one-tile entries on which EECBS performs poorly. Map (c) and Map (d) are maps similar to representative maps with two or more columns of empty space between each long obstacle component and short obstacle components on which EECBS performs well.
@@ Appendix C Maps with similar patterns
Figure 14 shows the easy maps of CBS, similar to the maps shown in Figure 3 in Section 4 in the main text. Figure 15 shows the challenging and easy maps of EECBS, similar to the maps shown in Figure 4 in Section 4 in the main text. Figure 16 shows the challenging and easy maps of PBS, similar to the maps shown in Figure 5 in Section 4 in the main text. Figure 17 shows the challenging and easy maps of PIBT, similar to the maps shown in Figure 6 in Section 4 in the main text. Figure 18 shows the challenging and easy maps of LTF, similar to the maps shown in Figure 7 in Section 4 in the main text.
Appendix D Validation Results
MAPF Algos Map Runtime Success Rate
CBS (fig. 3) Map (a) over 20s $0\%$
Map (b) 6.1s $78\%$
Map (c) over 20s $0\%$
Map (d) over 20s $0\%$
EECBS (fig. 4) Map (a) over 2.5s $97\%$
Map (b) 18.3s $5\%$
Map (c) over 20s $0\%$
Map (d) 0.08s $100\%$
PBS (fig. 5) Map (a) 3.4s $87\%$
Map (b) 3.6s $88\%$
Map (c) No solution $0\%$
Map (d) 19.1s $17\%$
PIBT (fig. 6) Map (a) \ $96\%$
Map (b) \ $95\%$
Map (c) \ $12\%$
Map (d) \ $53\%$
LTF (fig. 7) Map (a) \ $8\%$
Map (b) \ $27\%$
Map (c) \ $98\%$
Map (d) \ $30\%$
Table 1: Validation experiments results of generated maps in one-algorithm experiments. We run 200 MAPF instances in the generated maps and report the success rate and average runtime.
Map $T_{EECBS}$ $SR_{EECBS}$ $T_{PBS}$ $SR_{PBS}$
Map (a) 4.3s $95.5\%$ 12.5s $69.5\%$
Map (b) 9.8s $56.5\%$ 2.0s $96.5\%$
Map (c) 18.3s $14.5\%$ 13.7s $43.0\%$
Map (d) 1.1s $100\%$ 7.4s $77.5\%$
Table 2: Validation of observation by running maps with 200 different instances. $T$ is the average CPU runtime for success cases, and $SR$ is the success rate of 200 runs.
Map $SR_{PIBT}$ $SR_{LTF}$
Map (a) $70\%$ $56\%$
Map (b) $80\%$ $96\%$
Map (c) $52\%$ $80\%$
Map (d) $96\%$ $96\%$
Table 3: Validation of observation by running maps with 200 different instances. $SR$ is the success rate of 200 runs.
Table 1 shows the validation results of one-algorithm experiments in Section 4. Table 2 shows the validation results of two-algorithm experiments between EECBS and PBS in Section 5. Table 3 shows the validation results of two-algorithm experiments between PIBT and LTF in Section 5.